Let's say we have a mixture of h2 and o2. The mixture contained 6.7 hydrogen gas and 3.3 mols of oxygen gas. The mixture is in 300 l container at 273 Kelvin and total pressure of the gas mixture is 0.75atm. What is the partial pressure of each gas
1
Expert's answer
2020-05-01T14:30:20-0400

The partial pressure of each gas in a mixture of H2 and O2 is proportional to its mole fraction. As a result:

nx / nt = px / pt

where nx - number of moles of gas, nt - number of moles of a mixture, px - partial pressure of gas, pt - total pressure of mixture.

From here:

px = nx × pt / nt

Here, pt = 0.75 atm and nt = n(H2) + n(O2) = 6.7 mol + 3.3 mol = 10 mol

p(H2) = n(H2) × pt / nt = 6.7 mol × 0.75 atm / 10 mol = 0.50 atm

p(O2) = n(O2) × pt / nt = 3.3 mol × 0.75 atm / 10 mol = 0.25 atm


Answer: p(H2) = 0.50 atm, p(O2) = 0.25 atm